### v2

# Understanding In-context Learning of Addition via Activation Subspaces

Abstract
To perform in-context learning, language models must extract signals from individual few-shot examples, aggregate these into a learned prediction rule, and then apply this rule to new examples. How is this implemented in the forward pass of modern transformer models? To study this, we consider a structured family of few-shot learning tasks for which the true prediction rule is to add an integer $k$ to the input. We find that Llama-3-8B attains high accuracy on this task for a range of $k$, and localize its few-shot ability to just three attention heads via a novel optimization approach. We further show the extracted signals lie in a six-dimensional subspace, where four of the dimensions track the unit digit and the other two dimensions track overall magnitude. We finally examine how these heads extract information from individual few-shot examples, identifying a self-correction mechanism in which mistakes from earlier examples are suppressed by later examples. Our results demonstrate how tracking low-dimensional subspaces across a forward pass can provide insight into fine-grained computational structures.
footnotetext: Equal advising.\$*\$\$*\$footnotetext: Correspondence to: Xinyan Hu <xinyanhu@berkeley.edu>.
1 Introduction
Large language models (LLMs) have demonstrated remarkable in-context learning (ICL) capabilities, where they can adapt to various new tasks through few-shot demonstrations in the prompt. Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; Akyürek et al., 2023; von Oswald et al., 2023) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024). However, little is known about the fine-grained computational structure of in-context learning for large models.
To address this, we perform a detailed analysis of ICL for a structured family of tasks: learning to add a constant number (“add-k”). For these tasks, models are given an input sequence “$x_{1}\rightarrow y_{1}\#...\#x_{n}\rightarrow y_{n}\#x_{q}\rightarrow$” where $y_{i}=x_{i}+k$ for all $i$. This task is mathematically well-defined and tractable for an in-depth mechanistic analysis, while exhibiting the core challenge of extracting and applying a pattern from ICL examples. Moreover, modern language models are empirically capable of solving it (they correctly predict $y_{q}=x_{q}+k$).
Figure 1: In this paper, we study the activation space of attention heads in Llama-3-8B during in-context learning (ICL) for addition tasks. Our main findings: (1) out of $1024$ attention heads in Llama-3-8B, only three are important for these tasks (§4); (2) each head encodes the task information in a six-dimensional subspace (§4.1); (3) a six-dimensional subspace further decomposes into a four-dimensional subspace capturing the unit digit of the task information and a two-dimensional subspace capturing its ten digit (§4.3); and (4) task information is mostly extracted from a subset of tokens (§5). The figure illustrates these findings using an example add-k prompt.
To understand the mechanism behind in-context add-k, we study it in Llama-3-8B. We first discover that only a small number of model components are responsible for add-k: out of 1024 attention heads, only three do most of the work (§3). We find the task information (i.e., the number $k$) for each head is encoded in a six-dimensional subspace; zooming into each head, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4). We also examine how this head aggregates information from individual in-context examples (§5).
To discover the attention heads that are responsible for add-k, we introduce a novel approach inspired by previous work on function vectors (Todd et al., 2024), which are vectors that imitate ICL when they are patched into the residual stream (§2.2). We formulate an optimization method that identifies $33$ important heads whose outputs yield good ICL performance (§3.1). Through further ablation, we find that all but 3 of these 33 heads can be replaced with their mean value without significantly decreasing the performance (§3.2).
Since these three heads are responsible for computing add-k, they must extract signal about $k$ from individual examples and then represent $k$ in the residual stream. To understand how $k$ is represented, zooming into each head, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1). Furthermore, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions, and a two-dimensional subspace tracking the coarse-grained magnitude (§4.2-4.3). This offers insight on how LLMs internally represent addition, aligning with recent findings showing that (non-ICL) addition is encoded via trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025).
Finally, we examine how this head aggregates information from individual in-context examples. We find that the information is mostly extracted from the $y_{i}$ token positions, and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2). Moreover, we identify a self-correction mechanism: if the signal extracted from one token has an error, later tokens are likely to write in the opposite direction of the error (§5.3). This means that ICL goes beyond simple averaging of inputs and has stateful dynamics.
Our findings suggest that ICL tasks can be implemented through specialized activation subspaces from a small number of attention heads that extract, represent and aggregate task-related information in interpretable ways. We introduce a novel optimization method to identify important attention heads and demonstrate how a focused analysis of a simple ICL task can reveal delicate latent space structures and sophisticated computational strategies in a large language model.
2 Preliminaries
2.1 Model and Task
We begin by specifying the model and the task studied in this paper.
Model. We study Llama3-8B-Instruct, which has 32 layers and 32 attention heads per layer, and a residual dimension of 4096. We denote each head as a tuple (layer index, head index), where both indices range from 0 to 31.
Task.
For a constant $k$, the add-k task is to add $k$ to a given input integer $x$ to predict $y=x+k$. In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$” (where each $y_{i}=x_{i}+k$). The examples are concatenated in the prompt using the separator “##\##”, followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt. A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e. $y_{i}-x_{i}$ in the prompts) but not the input domain, in contrast to most previous studies (Todd et al., 2024; Hendel et al., 2023).
We construct data for the task by varying $x\in[1,100]$ and $k\in[1,30]$ (thus $y\in[2,130]$) since Llama3-8B is empirically capable of solving the addition task in this range. We consider the following three types of prompts:
  1. 1.
    five-shot ICL prompt, where all 5 demonstration examples satisfy $y_{i}=x_{i}+k$ for a fixed $k\in[1,30]$. We also call this add-k ICL.
  2. 2.
    mixed-$k$ ICL prompt, where the five demonstration examples are $y_{i}=x_{i}+k_{i}$ for five possibly different $k_{i}$ values.
  3. 3.
    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$” for some $x_{q}$.
2.2 Activation Patching and Function Vectors
Next, we briefly introduce activation patching, a common interpretability technique that will be used throughout the paper, and the function vector, a construction that we will use to identify important heads for our ICL tasks.
Activation patching.
Activation patching is performed by taking the activations of a model component when the model is run on one prompt, then patching in these activations when the model is run on a different prompt. Patching can either replace the model’s base activations or add to them; we will primarily consider the latter.
Specifically, if $z_{l}$ is the original value of the residual stream at layer $l$ at the final token position $\to$”, we patch in the replacement $z_{l}+v$, where $v$ is constructed from activations on a different input prompt. We choose the layer $l$ at one third of the network’s depth and construct $v$ from “function vector” heads, following Todd et al. (2024), as described next.
Function vectors.
Function vectors $v_{k}$ are vectors constructed from the outputs of selected attention heads, designed so that adding $v_{k}$ to the residual stream of a zero-shot prompt approximates the effect of the five-shot add-k task. For example, $v_{k}$ might be the average output of one or more attention heads across a set of five-shot add-$k$ examples. We define the intervention accuracy of $v_{k}$ as the average accuracy obtained on zero-shot prompts when $v_{k}$ is added to the residual stream across all tasks $k$. This metric captures how effectively $v_{k}$ encodes task-specific information about $k$. For comparison, we define clean accuracy as the accuracy on five-shot prompts without any intervention, also averaged across all tasks $k$.
In more detail, consider an attention head $h$: let $h(p)$ denote the output of head $h$ on prompt $p$ at the last token position, and define $h_{k}$ as the average of $h(p)$ across all five-shot add-k prompts. [^1] Similar to Todd et al. (2024), we approximate this average using $100$ randomly generated five-shot add-k prompts, where each prompt contains random demonstration examples $x_{i}$ and the final query input $x_{q}$ is chosen exactly once from each integer in $[1,100]$. Todd et al. (2024) identified a subset $\mathcal{H}$ of attention heads (for a different set of tasks) such that the vector $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ has high intervention accuracy—that is, adding it to the residual stream of zero-shot prompts effectively recovers few-shot task performance.
Building on this framework, we consider multiple ways to construct $v_{k}$: (1) the task-specific mean $h_{k}$ over the add-k task (as described above); (2) the overall mean $\bar{h}$ (i.e., average across all $k$), which removes task-specific information about $k$; or (3) the specific value $h(p)$ on an individual prompt $p$. Throughout the paper, we call $\bar{h}$ the mean-ablation of head $h$ and call $h_{k}$ the head vector of $h$. Beyond the three choices above, we sometimes project a head’s output onto a lower-dimensional subspace or scale it by a coefficient. Unlike Todd et al. (2024), we recover a different set of heads through solving a novel global optimization problem (§3.1) and systematic ablation studies (§3.2), which achieves higher performance (Figure 3).
Figure 2: The chain of localization in §3 and §4. We first identify $33$ significant attention heads (out of 1024) via a global optimization method (§3.1), then narrow down to 3 main heads while mean-ablating the remaining 30 heads (§3.2). We next study the structure of the representations of each main head by localizing it to a six-dimensional subspace (§4.1), and decompose it into a four-dimensional subspace encoding the unit digit and a two-dimensional subspace encoding the tens digit (§4.3).
3 Identifying Three Aggregator Heads
To understand the mechanism behind the add-k ICL task, we first need to find out what model components are responsible for performing it. In this section, we find that three heads do most of the work for add-k. We find these heads by first solving an optimization problem to identify $33$ significant heads out of 10241024102410243.1), and then further narrowing down to three main heads via systematic ablations (§3.2). This process is illustrated in the first row of Figure 2.
3.1 Identifying Significant Heads via Sparse Optimization
We will search for a set of heads that store the information for the add-k task, in the sense that their output activations yield good function vectors for add-k2.2).
Formulating the sparse optimization.
More formally, define $v_{k}(c)=\sum_{h}c_{h}\cdot h_{k}$, the sum of head outputs weighted by $c$, where $h$ goes over all $1024$ heads in the model. (Recall that $h_{k}$ is the average output of head $h$ averaged across a large dataset of five-shot add-k prompts.) We will search for a sparse coefficient vector $c\in[0,1]^{1024}$ such that adding $v_{k}(c)$ to the residual stream of a zero-shot prompt achieves high accuracy on the add-k task.
Let $\ell(x_{q},y_{q};v)$ be the cross-entropy loss between $y_{q}$ and the model output when intervening the vector $v$ onto the input “$x_{q}\to$”(i.e. replacing $z_{l}$ with $z_{l}+v$ in the forward pass on “$x_{q}\to$”, where $z_{l}$ is the layer-$l$ residual stream at the last token). We optimize $c$ with respect to the loss
$\mathcal{L}(c)=\mathbb{E}_{k\in[30]}\mathbb{E}_{x_{q}\in[100]}[\ell(x_{q},x_{q}+k;v_{k}(c))]+\lambda\|c\|_{1},$ (1)
where the regularization term with weight $\lambda$ promotes sparsity.
(a) Coefficients of heads.
(b) Average indirect effects of heads (Todd et al., 2024).
Figure 3: Comparison of significant heads identified by our optimized coefficients (left) and by average indirect effects (AIE) from the previous method (Todd et al., 2024) (right). Colors indicate the magnitude of each head’s importance (coefficients or AIE). The top $33$ heads identified by both methods are highlighted with frames ($13$ heads common across both methods in red and other $20$ heads in blue). Our identified heads yield an intervention accuracy of $0.85$, compared to the previous method’s accuracy of $0.31$. Both methods select heads from similar layers, but our optimization approach is significantly more effective.
Training details. We randomly select $25$ add-k tasks from our total $30$ tasks for training and in-distribution testing, and use the remaining five tasks only for out-of-distribution testing. For each task add-k, we generate $100$ zero-shot prompts “$x_{q}\rightarrow$”, where $x_{q}$ ranges over all integers from $[1,100]$ and the target output is $x_{q}+k$, yielding one data point for each (prompt, task) pair. We randomly split the data points of the $25$ tasks into training, validation, and test sets in proportions of $0.7$, $0.15$, and $0.15$, respectively. We use AdamW with learning rate $0.01$ and batch size $128$. We set the regularization rate $\lambda$ as $0.05$, which promotes sparsity while incurring little loss in accuracy. During training, we clip the coefficients $c$ back to $[0,1]$ if they go out of range after each gradient step.
Results. We get coefficients that achieves high intervention accuracy. In particular, at the final epoch, the average accuracies of intervening on test data for the $25$ tasks and the five tasks ($0.83$ and $0.87$) are both close to the average accuracies of doing no intervention ($0.89$ and $0.92$). To identify the important heads for the tasks, we plot the values of the coefficients in the final epoch for each layer and head index (Figure 3(a)). We find $33$ heads have coefficients $c_{h}$ greater than $0.2$, most ($21$) of which are one. In contrast, the other ($991$) coefficients are all smaller than $0.01$, most ($889$) of which are zero. We call the $33$ heads significant heads and denote the set of them as $\mathcal{H}_{\text{sig}}$.
Comparison to previous approach. We compare our optimization approach with the previous method from Todd et al. (2024) for identifying important heads, which selects heads based on average indirect effect (AIE). To perform a fair comparison, we construct our function vector by summing the outputs of our selected heads directly without weighting by coefficients, matching their methodology. Using this construction, we achieve an intervention accuracy of $0.85$, close to the clean accuracy of $0.87$, indicating that our $33$ heads captures most of the necessary information for the add-k task. In contrast, selecting the top 33 heads according to AIE yields a much lower intervention accuracy of $0.31$. [^2] Todd et al. select the top 10 heads in their work, but this yields an even lower accuracy of $0.05$ in our setting. In Figure 3, we visualize the coefficients and AIE values of heads from both methods and highlight the $13$ heads that are common to both approaches.
3.2 Further Refinement via Ablations
We suspect many heads are primarily responsible for storing formatting information (such as ensuring the output appears as a number) rather than encoding information about $k$ itself. Intuitively, while we require the overall signal transmitted by these heads, we do not need any information about the specific value of $k$. To test this hypothesis, we perform mean-ablations: replacing each task-specific signal $h_{k}$ with the overall mean $\overline{h}$ across all values of $k$2.2).
Specifically, we conduct mean-ablations over subsets of the 33 significant heads and measure the resulting intervention accuracy of the corresponding function vectors. Formally, when ablating a subset $\mathcal{H}_{0}$, the resulting function vector is given by
$v=\sum_{h\in\mathcal{H}_{0}}\overline{h}+\sum_{h\in\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}}h_{k}.$ (2)
Here, heads in $\mathcal{H}_{0}$ contribute only their overall mean signals, while heads in $\mathcal{H}_{\text{sig}}\setminus\mathcal{H}_{0}$ retain their task-specific information.
Focusing on two layers via layer-wise ablation. To efficiently narrow down the important heads, we first perform mean-ablations at the level of layers. From Figure 3(a), we observe that the significant heads $\mathcal{H}_{\text{sig}}$ are concentrated primarily in the middle and late layers. We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query. After trying different sets of layers, we found that mean-ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while mean-ablating any other combination of layers causes negligible drops in accuracy (Appendix A, Table 3). After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
Identifying three final heads via head-ablation. To understand the individual contributions of each head within layers 13 and 15, we perform mean-ablations at the level of individual heads. We first assess the intervention accuracy when retaining only the output of a single head while mean-ablating all other significant heads; however, this generally results in low accuracy. We hypothesize that the output magnitude of a single head is too small to significantly influence the model output, even if it encodes task-relevant information. To amplify each head’s effect, we scale its output by a coefficient (e.g., 5). We find that three heads—head 1 = $(15,2)$, head 2 = $(15,1)$, and head 3 = $(13,6)$—achieve intervention accuracies close to the clean accuracy when appropriately scaled, while all other heads show much lower accuracies regardless of scaling (Table 1). This suggests that these three heads individually encode the task information much better than any others.
Head  Coefficient   Accuracy
No intervention  N/A   0.87
[…]
Any other head  Optimal   0.19
Table 1: Intervened accuracies for scaling up each single head’s output by an optimal coefficient. Each of the top three heads (in red) achieves much higher intervention accuracy compared to any other significant head in layer $13$ and $15$ (in blue).
Finally, to remove the need for scaling while maintaining high intervention accuracy, we sum the outputs of these three heads (each with a coefficient of one) and mean-ablate all others. We find that summing the top three, top two, and only the top head yields intervention accuracies of $0.79$, $0.61$, and $0.21$, respectively. The three heads are thus collectively sufficient for performing the add-k task.
Validating necessity of the three heads via ablating them in five-shot ICL. So far, we have studied these three heads mainly through their contribution to the function vector $v_{k}$. We next directly test their necessity in the original five-shot ICL setting, by ablating the outputs of these three heads when running the model on many random five-shot ICL prompts. Our experiment shows that mean-ablating these three heads in five-shot prompts yields an accuracy of $0.43$, sharply decreasing from the clean accuracy $0.87$ by half. For comparison, we mean-ablate $20$ random sets of three significant heads (other than head 1,2,3); their accuracies remain close to the clean accuracy: $95\%$ of them have accuracy at least $0.86$.
4 Characterizing the Aggregator Subspace
For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL examples. Our next goal is thus to understand how task information is represented in the activation space. Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
In this section, we dissect their activation spaces in three stages: (1) Localize a six–dimensional task subspace in each head via principal component analysis (PCA) (§4.1); (2) Rotate this subspace into orthogonal feature directions aligned with sinusoidal patterns across $k$4.2); (3) Decompose the six-dimensional space into a four-dimensional unit-digit subspace and a two-dimensional magnitude subspace that separately encode the units and tens of the answer (§4.3).
4.1 Localizing to Six-dimensional Subspace
To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Figure 4). We then check that the function vectors found earlier remain effective after projecting onto the subspace. Specifically, we replace each head vector $h_{k}$ with its projection onto the subspace $\tilde{h}_{k}$ to obtain a new function vector $\tilde{v}_{k}=\sum_{h\in\mathcal{H}}\tilde{h}_{k}$ (a variant of $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ in §2.2). We find that $\tilde{v}_{k}$ has intervention accuracy $0.76$, which is close to the intervention accuracy of $0.79$ before projection (§3.2). Thus, we confine our study to this concise six-dimensional subspace for each head. The rest of the section focuses on head 1 and we defer analogous results for head 2 and 3 to appendix B.
Figure 4: The explained variance ratio vs. number of PCs for each head. The first six PCs make up most of the explained variance ($97\%$) for each head.
Figure 5: Coordinates of head 1’s vectors (their inner products with a PC) for the first six PCs across different add-$k$ tasks. The coordinates in the first five PCs reveal periodic patterns.
4.2 Identifying Feature Directions Encoding Periodic Patterns
To understand how the six-dimensional subspace of a head represents the task information, we first examine the coordinates of the head vectors (their inner products with the PCs) as a function of $k$. This reveals partially periodic patterns in the first five components (Figure 5).
This motivates us to linearly transform the six PCs to find directions that encode pure periodic patterns. Mathematically, if we can find a linear transformation of the six PC-coordinate functions that fits trigonometric functions, then by applying the transformation on the PCs, we can obtain six directions whose coordinate functions encode the periodicity.
To find trigonometric functions to fit, we searched over different periods and phases and performed least squares regression. We found six trigonometric functions at periods $2$, $5$, $10$, $10$, $25$, and $50$ that could be expressed as functions of the top 6 PCs with low regression error (Figure 6). We apply the resulting linear transformation to the six PCs to obtain a new set of directions that encode these six pure periodic patterns, which we call feature directions.
Figure 6: The coordinate functions (inner product with feature directions) of head 1’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
4.3 Decomposing to Subspaces Encoding Subsignals
Leveraging the feature directions identified previously, we decompose the head activation subspace into lower-dimensional components that separately encode different subsignals relevant to the task—in this case, the units digit and tens digit.
By construction, the coordinate function of each feature directions (viewed as linear projections of $h_{k}$) is a periodic function of $k$. Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo $T$”. Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We verify these hypotheses through causal intervention. In Figure 7, we validate hypothesis (ii): we establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal. We include analogous experimental results for the two other hypotheses of head 1 in Appendix B.2.
Figure 7: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto the “unit subspace” (left) and out of the “unit subspace” (right). Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully account for the final-answer errors. This confirms that the unit subspace specifically encodes the unit-digit signal.
5 Signal Extractors of ICL Examples
Previously, we localized the model’s behavior to three heads and their corresponding six-dimensional subspaces, then examined how the model represents the task information ($k$ for add-k) inferred from the ICL examples in one subspace. Now, we analyze how the model extracts the task information from the ICL examples.
In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting task information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior. For simplicity, we focus on head 1 in this section.
5.1 Mathematical Observation: Tracing Subspace back to Previous Tokens
We begin with a mathematical observation that lets us trace the subspace at the final token back to corresponding subspaces at earlier token positions. Intuitively, a head’s output at the last token is a weighted sum of transformed residual streams from the previous tokens, with the weights given by the attention scores. Thus, the signal extracted from previous tokens is the transformed residual stream at that token.
Formally, a head $h$’s output at the last token of a prompt $p$ can be written as
$h(p)=\sum_{t\in p}\alpha_{t}\cdot O_{h}V_{h}\cdot z_{t},$ (3)
where $\alpha_{t}$ is the attention score from the last token to each token $t$, satisfying $\sum_{t\in p}\alpha_{t}=1$, $z_{t}$ is the residual stream input to the head $h$ at token $t$, $V_{h}$ is the value matrix, and $O_{h}$ is the output matrix mapping from head-dimensional space to model-dimensional space.
Let $W_{h}$ denote the projection matrix onto the six-dimensional subspace for head $h$. Then the projected signal at the final token, $W_{h}\cdot h(p)$, can be decomposed into contributions from previous tokens as $W_{h}\cdot\alpha_{t}O_{h}V_{h}z_{t}$, each lying in the transformed head subspace defined by the matrix $W_{h}\cdot O_{h}V_{h}$. In the following subsection, we analyze the magnitudes and directions of these signals, and study how signals from different examples interact.
5.2 Signal Extractor for each Example
Figure 8: For a mixed-$k$ ICL prompt, the strength of the task-signal contribution of each previous token $t$ to the final token, $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$. Decomposing it into two parts: (1) the norm of the extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$ (top), and (2) the attention score from the final token $\alpha_{t}$ (bottom), both parts consistently peak at the tokens $t=y_{i}$ (in bold).
To understand how examples contribute to model generation at the last token, we identify which tokens contribute the most, then examine what information they provide. By the analysis above, the task-signal contribution of each previous token to the final token through the head $h$ is $\alpha_{t}W_{h}O_{h}V_{h}z_{t}$. This can be decomposed into two parts: (1) extracted information: $W_{h}O_{h}V_{h}z_{t}$, the residual stream input projected into the relevant subspace; and (2) aggregation weight: $\alpha_{t}$, the attention score of the final token to the previous token. We plot the norms of the extracted information and the aggregation weights for a random mixed-$k$ ICL prompt in Figure 8. Both the strength of the extracted information and the aggregation weights peak at $y_{i}$ tokens.
We next examine what specific information is extracted from each of these tokens. To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},\tilde{h}_{k}\rangle$ between the extracted information and the head vector (projected onto the subspace and normalized to have unit norm) for each task $k$. In Figure 9, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$. We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector (projected onto the subspace and normalized to have unit norm) $\tilde{h}_{k}$. For each example $x_{i}\rightarrow y_{i}$, the inner product peaks for $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
5.3 Signal Correlation among Examples
Having studied the signal extracted from each individual example in the previous subsection, we next study how signals from different examples interact to execute the ICL task. To do so, we compute the correlation between the extracted signals from different demonstration examples: for each $y_{i}$ token, we first compute the inner product between the residual stream input to head 1, $z_{t}$, and the corresponding task vector $h_{k}$, where $k=y_{i}-x_{i}$. Then, we compute the correlation of these measures across each pair of five positions over $100$ add-k prompts, yielding $\binom{5}{2}$ correlation values per task.
To analyze the correlation, we sum the negative correlation values and positive correlation values respectively for each task, and calculate the various statistics (max, average and min) over all tasks (Table 2). The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, indicating that the signals from any two demonstration examples are mostly negatively correlated. This suggests a self-correction mechanism: intuitively, when the head extracts a noisy signal from one example, signals from subsequent examples are more likely to correct the error, thereby stabilizing the final representation.
Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlations
Average $-2.01$ $0.27$
Min $-1.40$ $0.07$
Max $-2.34$ $0.54$
Table 2: The statistics (min, max, and average) of the absolute values of negative correlation sum and positive correlation sum respectively over the $30$ tasks. The negative correlation sum is significantly higher than the positive correlation sum for all three statistics, which indicates that the signals from any two demonstration examples are mostly negatively correlated, suggesting a self-correction mechanism.
6 Related Work
Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Olah et al., 2020; Wang et al., 2022; Wu et al., 2023; Cunningham et al., 2023; Feng & Steinhardt, 2023; Gould et al., 2023; Park et al., 2024). Below we discuss three particular threads that are most relevant to this paper.
Interpreting arithmetic tasks.
A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Hanna et al., 2023; Nikankin et al., 2024; Maltoni & Ferrara, 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025). Zhou et al. (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers. Different from them, our work is the first to analyze addition for in-context learning in LLM, and interestingly we find similar representation structures to Zhou et al. (2024) and Kantamneni & Tegmark (2025).
Interpreting in-context learning.
Researchers have constructed detailed models of in-context learning (ICL) for small transformer models in standard supervised learning problems such as linear regression (Garg et al., 2022; Akyürek et al., 2023; Zhang et al., 2023; Li et al., 2023; Wu et al., 2024), as well as more complex settings (von Oswald et al., 2023; Bai et al., 2023; Bietti et al., 2023; Reddy, 2023; Guo et al., 2023; Nichani et al., 2024). For large pretrained models, there exist coarser-grained treatments attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023). Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads. Motivated by this, we study function vector heads in details for a family of ICL tasks, introducing a novel optimization method, which achieves better performance than the method in Todd et al. (2024). Another difference from Todd et al. (2024) is that our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.
Causal analysis.
There has been a line of research that proposes methods to understand the causal influence of model components on model behavior, such as by probing (Conneau et al., 2018; Hewitt & Manning, 2019; Clark et al., 2019). Our methodological approach follows recent developments in revealing causal effects of model components by interventions on internal states of models (Vig et al., 2020; Geiger et al., 2021). In particular, we draw inspiration from causal mediation analysis used in Todd et al. (2024), activation patching (Meng et al., 2022), and causal scrubbing (Chan et al., 2022).
7 Discussion
In this work, we provided a detailed mechanistic analysis of in-context learning for addition tasks in Llama-3-8B. We found that a small number of attention heads operating in low-dimensional subspaces can extract, represent, and aggregate ICL task information in structured and interpretable ways. We performed our analysis in five steps:
  1. 1.
    Use sparse optimization to identify important attention heads whose outputs construct effective function vectors for ICL tasks (§3.1).
  2. 2.
    Localize task information to a smaller subset of heads via ablations (§3.2).
  3. 3.
    Further localize to low-dimensional subspaces via PCA on each remaining head (§4.1).
  4. 4.
    Examine this subspace qualitatively, which uncovered periodic patterns in the activation space (§4.2) that decomposed into interpretable subspaces encoding unit-digit and magnitude information (§4.3).
  5. 5.
    Exploit algebraic structure in the transformer to connect “aggregation” subspaces at the final token position with “extraction” subspaces at the earlier $y_{i}$ tokens (§5).
This same methodology (identify important heads, restrict to relevant subspaces, and examine the remaining information qualitatively) could be extended to other models and tasks. Most steps in our methodology also scale easily: the sparse optimization is fully automatic. While mean ablation involved some qualitative judgment, we can fold both of these steps into a single optimization task that mean ablates some heads while fully removing others. PCA is also automatic. For the final step that involves a qualitative examination of the subspaces, future work could explore automating this step using AI systems.
There are a few other questions stemming from our work. First, the three important heads are capable of individually representing and extracting information, but Llama3-8B relies on all of them—how does it jointly aggregate information from these three heads to produce a final answer? Second, we observe that heads implement a “self-correction” mechanism indicated by the negative correlation between examples signals, but the exact mechanism of this self-correction remains to be discovered.
Our findings demonstrate how a focused study can reveal sophisticated latent structures and delicate computational strategies within a LLM. We hope this motivates more work to dive deep into the LLM inner structures and obtain more understanding of LLM latent spaces and computations.
Acknowledgments
We thank Lisa Dunlap, Alex Pan, Tony Lian, Anya Ji, Jiaxin Ge, Junyi Zhang, Hanlin Zhu and Shu Liu for their helpful feedback and suggestions. KY is supported by the Vitalik Buterin Ph.D. Fellowship in AI Existential Safety. MJ is supported by the European Union (ERC-2022-SYG-OCEAN-101071601). LC is supported by a Miller Research Fellowship. JS is supported by Open Philanthropy, Simons Foundation, and NSF.
Appendix A Additional table in §3.2
In §3.2, we did systematic ablation studies to narrow down the tasks to three main heads from $33$ significant heads. The first step of the ablation studies is layer-wise ablation, where we mean-ablate the significant heads in a subset of layers. We include the experimental results here, which narrow down to layer 13 and 15.
Layer Accuracy
No intervention 0.87
$[0,31]$ 0.85
$[0,15]$ 0.83
$[13,15]$ 0.83
$\{13,15\}$ 0.83
$\{14,15\}$ 0.69
$\{13,14\}$ 0.25
$\{15\}$ 0.71
$\{13\}$ 0.27
$\{14\}$ 0.03
$[0,31]\setminus\{13,15\}$ 0.05
Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers. We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.87$, while other combinations lead to substantial drops in accuracy, especially when mean-ablating layers $13$ and $15$ (in blue).
Appendix B Additional figures in §4
In §4.2 and §4.3, we focused on head 1 for simplicity of exposition. Here we include the analogous experimental results for head 2 and 3.
B.1 Additional figures for §4.2
In §4.2, we found six trigonometric functions that can be linearly fitted by the coordinate functions of head 1 (Figure 6). Here we show analogous experimental results for head 2 and 3.
Figure 10: The coordinate functions (inner product with feature directions) of head 2’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
Figure 11: The coordinate functions (inner product with feature directions) of head 3’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
B.2 Additional figures for §4.3
In §4.3, we showed experimental results validating the hypothesis (ii) for head 1 (Figure 7). Here we show analogous experimental results validating the hypotheses (i) and (iii) for head 1, as well as all three hypotheses for heads 2 and 3. Recall that the three hypotheses are as follows:
  1. (i)
    the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task;
  2. (ii)
    the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$;
  3. (iii)
    the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We first show the experimental results validating hypotheses (i) and (iii) for head 1. Projecting out of the parity direction doesn’t lead to high errors for the parity and the final answer across all tasks, which might be because parity is relatively easy to obtain (e.g., random choice leads to 0.5 accuracy).
(a) Head 1 projected onto parity direction
(b) Head 1 projected out of parity direction
Figure 12: Validation of hypothesis (i) for head 1.
(a) Head 1 projected onto magnitude subspace
(b) Head 1 projected out of magnitude subspace
Figure 13: Validation of hypothesis (iii) for head 1.
We then show analogous results for head 2.
(a) Head 2 projected onto parity direction
(b) Head 2 projected out of parity direction
Figure 14: Validation of hypothesis (i) for head 2.
(a) Head 2 projected onto unit subspace
(b) Head 2 projected out of unit subspace
Figure 15: Validation of hypothesis (ii) for head 2.
(a) Head 2 projected onto magnitude subspace
(b) Head 2 projected out of magnitude subspace
Figure 16: Validation of hypothesis (iii) for head 2.
Finally, we show analogous results for head 3. The evidence for head 3 is slightly weaker than the first two heads, which aligns with the fact that the intervened accuracy of head 3 is slightly weaker than the ones of the first two heads (Tabel 1).
(a) Head 3 projected onto parity direction
(b) Head 3 projected out of parity direction
Figure 17: Validation of hypothesis (i) for head 3.
(a) Head 3 projected onto unit subspace
(b) Head 3 projected out of unit subspace
Figure 18: Validation of hypothesis (ii) for head 3.
(a) Head 3 projected onto magnitude subspace
(b) Head 3 projected out of magnitude subspace
Figure 19: Validation of hypothesis (iii) for head 3.